OpenStudy (anonymous):

Which percent is equivalent to 5 4/5? 560% 580% 590%

OpenStudy (jdoe0001):

\(\bf {\color{red} a}\%\ =\cfrac{{\color{red} a}}{100}\\ \quad \\ 5\cfrac{4}{5}\implies \cfrac{5\cdot 5+4}{5}\qquad\qquad \cfrac{5\cdot 5+4}{5}\%\implies \cfrac{\frac{5\cdot 5+4}{5}}{100}\)